OpenStudy (anonymous):

not a question Art of Problem Solving Methods of Solving Diophantine Equations 1. Discriminant of quadratic equation

OpenStudy (anonymous):

If the diophantine equation given is a quadratic equation with 2 variables, the discriminant can be very useful to determine the possible values of the variables. If the variables are stated to be positive integer, it suggests that the discriminant must be a perfect square (or in rare cases, the square of a rational number). ---------------------------------------------------------------------------- Here’s a problem (\(\color\red{\text {by myself}}\)) to introducing this method : Problem: Find all pairs of positive integers \(\color\red{\text {(m,n)}}\) such that : \(\color\red{\ { m+n=m^2-3mn+2n^2}}\) We have to arrange the equation to appear as a quadratic equation in terms of \(\color\red{\text {m}}\) or \(\color\red{\text {n}}\) .i do it for \(\color\red{\text {m}}\) : Upon rearrangement we have\(\color\red{\ {m^2-(3n+1)m+2n^2-n=0 }}\). By taking discriminant, we obtain \( \)\(\color\red{\ { \Delta=n^2+10n+1}}\). So the discriminant must be a perfect square but we have: \( (n+1)^2< \Delta=n^2+10n+1<(n+5)^2\) Hence: \( n^2+10n+1=(n+2)^2 , (n+3)^2 \ and \ (n+4)^2\) from this we get \(\color\red{\text {n=2}}\) and putting \(\color\red{\text {n=2}}\) in original equation gives \(\color\red{\text {m=1}}\) so the only solution is \(\color\red{\text {(m,n)=(1,2)}}\) ----------------------------------------------------------------------------

OpenStudy (vishweshshrimali5):

@mukushla Its very nice method

OpenStudy (anonymous):

Excellent... Sorry to say but I am not getting how did you find n from the last equation..?? Can you explain it to me?? @mukushla

OpenStudy (anonymous):

Yeah I got it from (n+3)^2..

OpenStudy (anonymous):

(n+1)^2=n^2+2n+1<n^2+10n+1<n^2+10n+25=(n+5)^2 so we have 3 options for n^2+10n+1 to be a perfect square n^2+10n+1=(n+2)^2 --------> n=1/2 not a integer n^2+10n+1=(n+3)^2 --------> n=2 n^2+10n+1=(n+4)^2 --------> n=15/2 not a integer

mathslover (mathslover):

must say that : great tutorial .. \[\Huge{\text{Great}\mathbb{Job!}\textbf{Keep}\mathbb{It}\textbf{Up}}\]

OpenStudy (anonymous):

@mathslover tnx my friend...

mathslover (mathslover):

No thanks , u were the desrver

OpenStudy (anonymous):

I wish I was good at maths like you..I'm so dumb :(

OpenStudy (jiteshmeghwal9):

\[\LARGE{Great \space job \space Dude \space ;)}\]

OpenStudy (anonymous):

I wish there was a search function on this site :-D gj m8!

OpenStudy (ujjwal):

Superb!!

OpenStudy (anonymous):

great job! i used such things before,but never found it anywhere so clear!

OpenStudy (anonymous):

@quarkine happy to hear that

OpenStudy (anonymous):

i hope u give me permition to save this link... thanks

OpenStudy (anonymous):

sure...

OpenStudy (ikram002p):

nice :D